An air refrigerator works between the pressure limits of 1 bar and 5 bar. The temperature of air entering the compressor and expansion cylinder are 10oC and 25oC respectively. The expansion and compression follow the law pv1.3=constant. Find the following
1
Expert's answer
2020-06-05T05:36:43-0400

Here, "P_1= 1 bar" and "P_2=" 5 bar is given

"T_1=10^oC, T_2 = 25 ^oC"

It is given that it is following the following equation

"PV^{1.3}=C"


Here we can find different values as per requirement but it is not mention in question wat we have to find


Question is not complete